# Leveraging 3D Representation Alignment and RGB Pretrained Priors for LiDAR Scene Generation

### v1

@@ II Related works
Some LiDAR scan generation methods, such as UltraLiDAR [6] and OLiDM [28], directly take point clouds as input. To control for non-uniform and large point cloud sizes, these methods rely on encoding and pre-processing strategies such as voxelization or padding to create uniform point clouds. The second and most widely used modality, range images, requires no costly pre-processing, and builds on top of existing classic image architectures like UNet [29] or ViT [30]. Range images enable an easier fusion between modalities and the possibility to use large pretrained models [31].
First attempts to use diffusion on range images opted for pixel space diffusion [5, 21], which comes at the price of a costly generation process limiting the size of the network. More recent works, such as [23, 22], switch to a latent diffusion approach [10]. Conversely, R2Flow [20] argues that diffusion in pixel space is essential for high-fidelity results. To reduce inference time, it introduces a rectified flow model that lowers the number of required sampling steps. In contrast, we argue that latent diffusion can also deliver high-quality generation, provided the variational autoencoder is trained with an adversarial loss as shown in [22].
@@ IV-A Experimental Setup
Implementation Details. We use 3D self-supervised features from ScaLR [39], which ranks among the state-of-the-art self-supervised backbones for LiDAR. We implement our FM model using the base Scalable Interpolant Transformer [41] with a patch size of 2 (SiT-B/2). We train SiT-B/2 for 1M steps with a batch size of 256 using the Stable Diffusion VAE on ImageNet, following the official code and protocol of [14]. During the VAE alignment stage, we train our VAE from scratch for 800k steps with a batch size of 32. After alignment, we perform end-to-end training for 1M steps, also with a batch size of 32. We set $\lambda_{1}$ and $\lambda_{2}$ to 1.5 and 0.5, respectively.
@@ IV-B Comparison to state-of-the-art
For each method, we generate 10,000 samples and evaluate them against the full dataset of real scenes (train-val evaluation), as is common in generative modeling [20]. For completeness, we also report results evaluated against the validation set (val-only evaluation). For MMD, which requires computing the expensive Chamfer distance between all pairs of point clouds, we use 2,000 generated samples and evaluate them against 1,000 validation scenes, as in [23].
We present quantitative results in Table I and qualitative comparisons in Figure 5. For range-image metrics, R3DPA performs on par with the state of the art on FRID, while achieving a significant improvement on the FLD metric. For point-based metrics, which are particularly important given the ultimate goal of generating high-quality, realistic point clouds, our method surpasses previous approaches. On bird’s-eye-view metrics, R3DPA improves the JSD score while achieving performance comparable to the state of the art on MMD. Importantly, we achieve strong performance across both point and range metrics while other works show higher variance. We present further qualitative results from our model in Figure 6.
@@ IV-D Test-Time Conditioning with 3D Representations
Real scene Generated scene Real scene
Figure 8: Test-time conditioning with 3D representations for scene mixing. We condition on stitched features from two real scenes, mixing the left half of one scene with the right half of the other. Generated scene (middle) mixes layouts from two real scenes: the left part (purple) resembles the first real scene (left), while the right part (green) resembles the second real scene (right).
@@ V Conclusion
In this work, we introduced R3DPA, the first unconditional LiDAR point cloud generation method that integrates RGB image pretrained priors and self-supervised 3D representations. Our proposed training strategy transfers knowledge from large-scale image flow matching models, compensating for the limited scale of LiDAR datasets, while representation alignment with self-supervised 3D features further improves generation quality. Beyond synthesis, our feature-guided training enables controllable editing at inference, where we present first promising results for object inpainting and scene mixing. We believe this axis will inspire further research. Experiments on KITTI-360 show that R3DPA sets a new state of the art, opening new directions for bridging 2D and 3D representation learning in LiDAR generative modeling.
@@ ACKNOWLEDGMENT
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic.

### v2

# R3DPA: Leveraging 3D Representation Alignment and RGB Pretrained Priors for LiDAR Scene Generation

@@ II Related works
Some LiDAR scan generation methods, such as UltraLiDAR [6] and OLiDM [28], directly take point clouds as input. To control for non-uniform and large point cloud sizes, these methods rely on encoding and pre-processing strategies such as voxelization or padding to create uniform point clouds [29]. The second and most widely used modality, range images, requires no costly pre-processing, and builds on top of existing classic image architectures like UNet [30] or ViT [31]. Range images enable an easier fusion between modalities and the possibility to use large pretrained models [32].
First attempts to use diffusion on range images opted for pixel space diffusion [5, 21, 33], which comes at the price of a costly generation process limiting the size of the network. More recent works, such as [23, 22], switch to a latent diffusion approach [10]. Conversely, R2Flow [20] argues that diffusion in pixel space is essential for high-fidelity results. To reduce inference time, it introduces a rectified flow model that lowers the number of required sampling steps. In contrast, we argue that latent diffusion can also deliver high-quality generation, provided the variational autoencoder is trained with an adversarial loss as in [22].
@@ IV-A Experimental Setup
Implementation Details. We use 3D self-supervised features from ScaLR [41], which ranks among the state-of-the-art self-supervised backbones for LiDAR. We implement our FM model using the base Scalable Interpolant Transformer [43] with a patch size of 2 (SiT-B/2). We train SiT-B/2 for 1M steps with a batch size of 256 using the Stable Diffusion VAE on ImageNet, following the official code and protocol of [14]. During the VAE alignment stage, we train our VAE from scratch for 800k steps with a batch size of 32. After alignment, we perform end-to-end training for 1M steps, also with a batch size of 32. We set $\lambda_{1}$ and $\lambda_{2}$ to 1.5 and 0.5, respectively, as in [39].
@@ IV-B Comparison to state-of-the-art
For each method, we generate 10,000 samples and evaluate them against the full dataset of real scenes (train-val evaluation), as is common in generative modeling [20]. For completeness, we also report results evaluated against the validation set (val-only evaluation). For MMD, which requires computing the expensive Chamfer distance between all pairs of point clouds, we use 2,000 generated samples and evaluate them against 1,000 validation scenes, as in [23]. The VAE alignment and end-to-end training require 48h and 70h, respectively, using 8 NVIDIA A100 GPUs. R3DPA achieves an average inference speed of 0.6 s/scene, generating 10k samples in 1.45h on a single A100 with a batch size of 64.
We present quantitative results in Table I and qualitative comparisons in Figure 5. For range-image metrics, R3DPA performs on par with the state of the art on FRID, while achieving a significant improvement on the FLD metric. For point-based metrics, which are particularly important given the ultimate goal of generating high-quality, realistic point clouds, our method surpasses previous approaches. On bird’s-eye-view metrics, R3DPA improves the JSD score while achieving performance comparable to the state of the art on MMD. Importantly, we achieve strong performance across both point and range metrics while other works show higher variance. We suspect that [6]’s low score likely results from its voxel-based architecture. We present further qualitative results from our model in Figure 6.
@@ IV-D Test-Time Conditioning with 3D Representations
Real scene - I Generated scene Real scene - II
Figure 8: Test-time conditioning with 3D representations for scene mixing. We condition on stitched features from two real scenes, mixing the left half of one scene with the right half of the other. Generated scene (middle) mixes layouts from real scenes: the left part (purple) resembles the first real scene (left), while the right part (green) resembles the second real scene (right).
@@ V Conclusion
In this work, we introduced R3DPA, the first unconditional LiDAR point cloud generation method that integrates RGB image pretrained priors and self-supervised 3D representations. Our proposed training strategy transfers knowledge from large-scale image flow matching models, compensating for the limited scale of LiDAR datasets, while representation alignment with self-supervised 3D features further improves generation quality. Beyond synthesis, our feature-guided training enables controllable editing at inference, where we present first promising results for object inpainting and scene mixing. We believe this axis will inspire further research. Another promising next step involves leveraging these generations for downstream tasks such as LiDAR segmentation. Experiments on KITTI-360 show that R3DPA sets a new state of the art, opening new directions for bridging 2D and 3D representation learning in LiDAR generative modeling.
@@ ACKNOWLEDGMENT
We acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-REG-2024R02-234 access to Karolina, Czech Republic.